OpenStudy (anonymous):

Find the area of a circle with diameter 8 cm.

OpenStudy (anonymous):

200.96

OpenStudy (blockcolder):

Hint: What is the radius?

OpenStudy (anonymous):

oh sorry

OpenStudy (anonymous):

Do you know the formula?

OpenStudy (anonymous):

50.24 is the right answer

OpenStudy (anonymous):

(pie)r^2

OpenStudy (anonymous):

i was asking her Drake lol;P

OpenStudy (anonymous):

ya, i was just giving it to her

OpenStudy (anonymous):

the radius is 4

OpenStudy (callisto):

Area of circle = πr^2 , where r is the radius of the circle. r = d/2 , where d is the diameter of the circle. Now, you have d =8. r=d/2 = 8/2 =4 Area of circle = πr^2 = π(4)^2 = 16π cm^2 @DoomDude You shouldn't give answer only. You should also show how you get the answer so that the asker can learn and know what you have done to get the answer. Please read the Code of Conduct : http://openstudy.com/code-of-conduct